QUESTION 1 5 Find the Laplace transform of the function f(t) t, 0<t<1 1, t > 1
Answer #1


f(t)=st ostsi ] t>,1 Ste F (S) = -st dt +flest dt 6 Elt =ltest + - S remote) +($) +1 2 2 + 0 + 1 / 2 1 =-e-s -S1 as te -S = e